Determine vector A x C if A = 5i - 4j and B = 5i + 2j


1
Expert's answer
2022-03-24T16:36:42-0400

Answer

Vectors

A = 5i - 4j

B = 5i + 2j

Then

"A\times B=\begin{vmatrix}
 i & j & k\\
 5 & -4 & 0\\5 &2&0
\end{vmatrix}"


="i(0) -j(0) +k(10+20) \\=30k"